OpenStudy (anonymous):

Solve the equation: 5x^2=19x+4

OpenStudy (anonymous):

Use algebra to move everything to one side of the equation.. Then use the quadratic formula to solve it.

OpenStudy (anonymous):

5x^2 - 19x -4 =0 like he said or you can factor it (5x+1)(x-4)=0 x=-1/5, +4

OpenStudy (anonymous):

ty I wasn't sure if -1/5 was one